# SGS-SLAM: Semantic Gaussian Splatting for Neural Dense SLAM

### 02 Mar 2024 v3

@@ Abstract
Keywords SLAM  $\cdot$ 3D Reconstruction  $\cdot$ 3D Semantic Segmentation
@@
**footnotetext: These authors contributed equally to this workfootnotetext: Corresponding author
1 Introduction
Dense Visual Simultaneous Localization and Mapping (SLAM) is a crucial problem in the field of computer vision. It aims to reconstruct a dense 3D map in an unseen environment while simultaneously tracking the camera poses in a real-time manner. Traditional visual SLAM systems Davison et al. (2007); Newcombe et al. (2011); Salas-Moreno et al. (2013); Mur-Artal et al. (2015) stand out in sparse reconstruction using point clouds and voxels, but fall short in dense reconstruction. To extract dense geometric information for high-quality representation, learning-based SLAM methods Bloesch et al. (2018); Sucar et al. (2020) have gained wild attention. They demonstrate proficiency in generating decent 3D global maps meanwhile exhibiting robustness on noises and outliers. In addition, drawing inspiration from the advancements in the neural radiance field (NeRF) Mildenhall et al. (2021), NeRF-based SLAM approaches Sucar et al. (2021); Zhu et al. (2022); Kong et al. (2023); Zhang et al. (2023); Li et al. (2023a); Wang et al. (2023) have made further progress. They excel in producing accurate and high-fidelity global reconstruction by capturing dense photometric information through differentiable rendering.
@@ 1 Introduction
In this context, as opposite to NeRF-based neural representation, our exploration shifts towards the volumetric representation based on the 3D Gaussian Radiance Field Kerbl et al. (2023). This approach marks a significant shift and offers notable advantages in the scene representation.
@@
When compared with neural implicit semantic SLAM systems, such as DNS-SLAM Li et al. (2023a) and SNI-SLAM Zhu et al. (2023), our system demonstrates remarkable superiority in terms of rendering speed, reconstruction quality, and segmentation accuracy. Leveraging these benefits, our method enables precise editing and manipulation of specific scene elements while preserving the high fidelity of the overall rendering. Furthermore, using explicit spatial and semantic information to identify scene content can be instrumental in optimizing camera tracking. Particularly, we incorporate a two-level adjustment based on geometric and semantic criteria for keyframes selection. This process relies on recognizing objects that have been previously seen in the trajectory. Extensive experiments are conducted on both synthetic and real-world scene benchmarks. These experiments compare our method against both implicit NeRF-based approaches Yang et al. (2022); Zhu et al. (2022); Wang et al. (2023); Johari et al. (2023), and novel 3D-Gaussian-based methods Keetha et al. (2023), evaluating performance in mapping, tracking, and semantic segmentation.
@@ 2.1 Semantic SLAM
Semantic information is of great importance for SLAM systems Mur-Artal et al. (2015); He et al. (2023); Whelan et al. (2015); Qin et al. (2018), which is a crucial requirement for applications in robotics and VR or AR fields. Real-time dense semantic SLAM systems Salas-Moreno et al. (2013); Bloesch et al. (2018); Rosinol et al. (2020) integrate semantic information into 3D geometric representations. Traditional semantic SLAM systems rely on sparse 3D semantic expressions, such as voxel Hermans et al. (2014), point cloud Narita et al. (2019), and signed distance field Narita et al. (2019). These methods struggle with accurately interpreting complex environments due to limited semantic understanding. This results in a simplified categorization of environmental features, which may not capture the full range of objects and their relationships within a space. Moreover, these methods exhibit limitations regarding reconstruction speed, high-fidelity model acquisition, and memory usage.
@@ 2.2 Neural Implicit SLAM
Methods based on NeRF McCormac et al. (2018), which handle complex topological structures and differentiable scene representation methods, have garnered significant attention, leading to the development of neural implicit SLAM methods Chung et al. (2023); Li et al. (2023b, 2024); Zhou et al. (2024). iMAP Sucar et al. (2021) uses a single MLP for scene representation, which shows limitations in large-scale scenes. NICE-SLAM Zhu et al. (2022) uses pre-trained multiple MLPs for hierarchical scene representation. Co-SLAM Wang et al. (2023) combines pixel set-based keyframe tracking with one-blob encoding. Go-SLAM Zhang et al. (2023) uses Droid-SLAM as a front-end tracking system and multi-resolution hash encoding Müller et al. (2022) for mapping meanwhile integrated loop-closure detection and global optimization. However, these methods cannot utilize semantic information in maps. NIDS-SLAM Haghighi et al. (2023) leverages the mature front-end tracking of ORB-SLAM3 Campos et al. (2021) and Instant-NGP Müller et al. (2022) for mapping but does not optimize joint semantic features for 3D reconstruction. DNS-SLAM Li et al. (2023a) proposes a 2D semantic prior system that provides multi-view geometry constraints but does not optimize 3D reconstruction with semantic features. SNI-SLAM Zhu et al. (2023), a work parallel to ours, introduces semantic loss for geometric supervision but remains limited by the efficiency constraints of NeRF’s volume rendering.
@@
The outstanding performance and fast rasterization capabilities of 3D Gaussian Splatting Kerbl et al. (2023) enable higher efficiency and accuracy on scene reconstruction. However, existing 3DGS-based SLAM systems Yan et al. (2023); Keetha et al. (2023); Huang et al. (2023) lack traditional effective tracking optimization, limiting tracking accuracy and the ability to recognize semantic information in scenes. We fuse semantic features into geometry and appearance and integrate 3D semantic features during keyframe selection in the tracking process. This allows us to obtain more effective and higher-resolution scene segmentation results while maintaining real-time performance.
@@ 3 Method
SGS-SLAM is a Gaussian-based semantic visual SLAM system. section 3.1 introduces its multi-channel Gaussian representation for joint parameter optimization. Like previous SLAM techniques, our method can be split into two processes: tracking and mapping. The tracking process estimates the camera pose of each frame while keeping the scene parameters fixed. Mapping optimizes the scene representations based on the estimated camera pose. section 3.2 explains the breakdown steps in detail. In addition, section 3.3 presents scene manipulation as a case study for downstream tasks. fig. 1 shows an overview of our system.
@@ 3.1 Multi-Channel Gaussian Representation
The scene is represented using a Gaussian influence function $f(\cdot)$ on the map, For simplicity, these Gaussians are isotropic, as proposed in Keetha et al. (2023):
@@
In order to optimize the parameters of Gaussians to represent the scene, we need to render the Gaussians into 2D images in a differentiable manner. We use the render method from Luiten et al. (2024), providing extended functionality of rendering depth in colors. It works by splatting 3D Gaussians into the image plane via approximating the integral projection of the influence function $f(\cdot)$ along the depth dimension in pixel coordinates. The center of the Gaussian $\mu$, radius $r$, and depth $d$ (in camera coordinates) is splatted using the standard point rendering formula:
@@
where $K$ is the camera intrinsic matrix, $E_{t}$ is the extrinsic matrix capturing the rotation and translation of the camera at frame $t$, $l$ is the focal length. The influence of all Gaussians on this pixel can be combined by sorting the Gaussians in depth order and performing front-to-back volume rendering using the Max volume rendering formula Max (1995):
@@
where $d_{i}$ denotes the depth of each Gaussian. By setting $d_{i}=1$, we can calculate a silhouette, $Sil_{\rm pix}=D_{\rm pix}(d_{i}=1)$, which assists in determining whether a pixel is visible in the current view. This aspect of visibility is essential for camera pose estimation, as it relies on the current reconstructed map. Additionally, it is also employed in map reconstruction, where new Gaussians are introduced in pixels lacking sufficient information.
@@ 3.2.3 Map Reconstruction
$\mathcal{L}_{\rm mapping}=\mathcal{U}\sum_{\rm pix}\lambda_{D}|D_{\rm pix}^{GT}-D_{\rm pix}|+\lambda_{C}\mathcal{L}_{C}+\lambda_{S}\mathcal{L}_{S}$ (8)
where $\mathcal{L}_{C}$ and $\mathcal{L}_{S}$ are weighted SSIM loss Kerbl et al. (2023) with respect to appearance image and semantic image:
@@
$\lambda_{D}$, $\lambda_{C}$, $\lambda_{S}$, and $\alpha$ are predefined hyperparameters, and $\mathcal{U}$ is the uncertainty score defined in section 3.2.2.
Compared to existing NeRF-based approaches Zhu et al. (2022); Johari et al. (2023); Li et al. (2023a); Zhu et al. (2023) that necessitate complex model architectures and feature fusion strategies, SGS-SLAM adopts explicit Gaussian representation for mapping. This approach allows for direct gradient flow to each parameter, resulting in high rendering speeds and optimal reconstruction quality. Compared to recent Gaussian-based methods Keetha et al. (2023); Yan et al. (2023), SGS-SLAM incorporates geometric, appearance, and semantic features for multi-channel rendering. This enables the joint optimization of parameters across different channels, remarkably enhancing the efficiency and effectiveness of both mapping and segmentation processes.
@@ 4.1.1 Datasets
We evaluate our method on both synthetic and real-world datasets. To compare with other neural implicit SLAM methods, we evaluate synthetic scenes from Replica dataset Straub et al. (2019) and real-world scenes from ScanNet Dai et al. (2017a) and ScanNet++ Yeshwanth et al. (2023) datasets. The ground-truth camera pose and semantic map of Replica are offered from simulation, and the ground-truth camera pose of ScanNet is generated by BundleFusion Dai et al. (2017b). The ground-truth 2D semantic label is provided by the dataset.
@@ 4.1.3 Baselines
We compare the tracking and mapping with state-of-the-art methods iMap Sucar et al. (2021), NICE-SLAM Zhu et al. (2022), Co-SLAM Wang et al. (2023), ESLAM Johari et al. (2023), and SplaTAM Keetha et al. (2023). For semantic segmentation accuracy, we compare with NIDS-SLAM Haghighi et al. (2023), DNS-SLAM Li et al. (2023a), and SNI-SLAM Zhu et al. (2023).
@@ 4.2 Evaluation of Mapping and Localization
We show quantitative measures of reconstruction quality using the Replica dataset Straub et al. (2019) in table 1. Our method demonstrates state-of-the-art performance. Compared to other baseline methods, our approach attains notably superior outcomes, outperforming them by a margin of 10dB in PSNR.
Figure 2: Qualitative comparison of our method and the baselines for reconstruction across three scenes from the Replica Dataset Straub et al. (2019), with key details accentuated using colorful boxes. The results demonstrate that our method delivers more high-fidelity and robust reconstructions, particularly by capturing more detailed features of the objects in the scene.
In fig. 2, we present the reconstruction results of three chosen scenes, where regions of interest are accentuated with boxes in various colors. Our method exhibits high-fidelity reconstruction outcomes. Specifically, for small, intricately textured objects like a clock, socket, books on a tea table, and a lamp, our approach shows remarkable accuracy over NeRF-based methods. This is because Gaussians are capable of representing objects with complex textures and surfaces. Furthermore, NeRF-based methods often struggle with the over-smoothing issue, resulting in blurred edges on objects. In contrast, by utilizing an explicit Gaussian representation, SGS-SLAM precisely captures objects with clear edges, irrespective of their sizes. Compared with SplaTAM, which is also a Gaussian-based model, our approach utilizes semantic information for discerning object categories, recognizing visual appearance to determine texture, and applying geometric constraints to preserve accurate shapes. This combination enables our method to achieve thorough modeling of both objects and their surrounding environment. The combination of these constraints allows SGS-SLAM to capture fine-grained details of objects, offering high-fidelity and accurate reconstruction.
Table 1: Quantitative comparison of our method and the baselines in training view rendering on the Replica dataset Straub et al. (2019). Our method demonstrates SOTA performances on all three metrics.
Methods Metrics Avg. Room0 Room1 Room2 Office0 Office1 Office2 Office3 Office4
@@
[…]
Table 2: Quantitative comparison in terms of Depth L1, ATE, memory usage, and FPS between our method and the baselines on the Replica dataset Straub et al. (2019). The values represent the average outcomes across eight scenes. The results of baselines are retrieved from Zhu et al. (2023). Our method remarkably outperforms the baselines at Depth and ATE evaluations, and performs fairly on FPS metrics. Note that for parameter size, we only count the parameters of the model.
[…]
iMAP 4.645 3.118 4.153 9.92 2.23 1.82
[…]
table 2 displays the tracking evaluation results on the Replica dataset Straub et al. (2019). Our method excels in achieving the highest level of depth L1 loss (cm) and minimal ATE error, surpassing baseline methods by 70% in terms of depth loss and 34% in terms of ATE RMSE (cm). This exceptional performance can be attributed to our precise scene reconstruction, which provides finely-detailed rendering results. The high-quality rendering, in turn, contributes to accurate camera pose estimation based on the established map by preventing incorrect geometric reconstruction, which could otherwise result in inaccurate tracking outcomes. Additionally, utilizing features from different channels of Gaussians, such as geometry, appearance, and semantic information, provides multiple levels of supervision, resulting in a more robust and accurate tracking capability.

### 13 Mar 2024 v4

@@ Abstract
Keywords:
SLAM 3D Reconstruction 3D Semantics 3D Segmentation
**footnotetext: These authors contributed equally to this workfootnotetext: Corresponding author  footnotetext: The demo video can be found at https://youtu.be/y83yw1E-oUo.
1 Introduction
@@
Dense Visual Simultaneous Localization and Mapping (SLAM) is a crucial problem in the field of computer vision. It aims to reconstruct a dense 3D map in an unseen environment while simultaneously tracking the camera poses in a real-time manner. Traditional visual SLAM systems [6, 29, 32, 27, 39, 9] stand out in sparse reconstruction using point clouds and voxels, but fall short in dense reconstruction. To extract dense geometric information for high-quality representation, learning-based SLAM methods [1, 36] have gained wild attention. They demonstrate proficiency in generating decent 3D global maps meanwhile exhibiting robustness on noises and outliers. In addition, drawing inspiration from the advancements in the neural radiance field (NeRF) [25], NeRF-based SLAM approaches [35, 47, 18, 44, 19, 37, 8] have made further progress. They excel in producing accurate and high-fidelity global reconstruction by capturing dense photometric information through differentiable rendering.
@@
In this context, as opposite to NeRF-based neural representation, our exploration shifts towards the volumetric representation based on the 3D Gaussian Radiance Field [17]. This approach marks a significant shift and offers notable advantages in the scene representation.
@@
When compared with neural implicit semantic SLAM systems, such as DNS-SLAM [19] and SNI-SLAM [46], our system demonstrates remarkable superiority in terms of rendering speed, reconstruction quality, and segmentation accuracy. Leveraging these benefits, our method enables precise editing and manipulation of specific scene elements while preserving the high fidelity of the overall rendering. Furthermore, using explicit spatial and semantic information to identify scene content can be instrumental in optimizing camera tracking. Particularly, we incorporate a two-level adjustment based on geometric and semantic criteria for keyframes selection. This process relies on recognizing objects that have been previously seen in the trajectory. Extensive experiments are conducted on both synthetic and real-world scene benchmarks. These experiments compare our method against both implicit NeRF-based approaches [41, 47, 37, 15], and novel 3D-Gaussian-based methods [16], evaluating performance in mapping, tracking, and semantic segmentation.
@@ 2.1 Semantic SLAM
Semantic information is of great importance for SLAM systems [27, 12, 38, 30], which is a crucial requirement for applications in robotics and VR or AR fields. Real-time dense semantic SLAM systems [32, 1, 31] integrate semantic information into 3D geometric representations. Traditional semantic SLAM systems rely on sparse 3D semantic expressions, such as voxel [13], point cloud [28], and signed distance field [28]. These methods struggle with accurately interpreting complex environments due to limited semantic understanding. This results in a simplified categorization of environmental features, which may not capture the full range of objects and their relationships within a space. Moreover, these methods exhibit limitations regarding reconstruction speed, high-fidelity model acquisition, and memory usage.
@@ 2.2 Neural Implicit SLAM
Methods based on NeRF [24], which handle complex topological structures and differentiable scene representation methods, have garnered significant attention, leading to the development of neural implicit SLAM methods [3, 21, 7, 20, 45]. iMAP [35] uses a single MLP for scene representation, which shows limitations in large-scale scenes. NICE-SLAM [47] uses pre-trained multiple MLPs for hierarchical scene representation. Co-SLAM [37] combines pixel set-based keyframe tracking with one-blob encoding. Go-SLAM [44] uses Droid-SLAM as a front-end tracking system and multi-resolution hash encoding [26] for mapping meanwhile integrated loop-closure detection and global optimization. However, these methods cannot utilize semantic information in maps. NIDS-SLAM [11] leverages the mature front-end tracking of ORB-SLAM3 [2] and Instant-NGP [26] for mapping but does not optimize joint semantic features for 3D reconstruction. DNS-SLAM [19] proposes a 2D semantic prior system that provides multi-view geometry constraints but does not optimize 3D reconstruction with semantic features. SNI-SLAM [46], a work parallel to ours, introduces semantic loss for geometric supervision but remains limited by the efficiency constraints of NeRF’s volume rendering.
@@
The outstanding performance and fast rasterization capabilities of 3D Gaussian Splatting [17] enable higher efficiency and accuracy on scene reconstruction. However, existing 3D-Gaussian-based SLAM systems [40, 16, 14] lack traditional effective tracking optimization, limiting tracking accuracy and the ability to recognize semantic information in scenes. We fuse semantic features into geometry and appearance and integrate 3D semantic features during keyframe selection in the tracking process. This allows us to obtain more effective and higher-resolution scene segmentation results while maintaining real-time performance.
@@ 3 Method
SGS-SLAM is a Gaussian-based semantic visual SLAM system. Sec. 3.1 introduces its multi-channel Gaussian representation for joint parameter optimization. Like previous SLAM techniques, our method can be split into two processes: tracking and mapping. The tracking process estimates the camera pose of each frame while keeping the scene parameters fixed. Mapping optimizes the scene representations based on the estimated camera pose. Sec. 3.2 explains the breakdown steps in detail. In addition, Sec. 3.3 presents scene manipulation as a case study for downstream tasks. Fig. 1 shows an overview of our system.
@@ 3.1 Multi-Channel Gaussian Representation
The scene is represented using a Gaussian influence function $f(\cdot)$ on the map, For simplicity, these Gaussians are isotropic, as proposed in [16]:
@@
In order to optimize the parameters of Gaussians to represent the scene, we need to render the Gaussians into 2D images in a differentiable manner. We use the render method from [22], providing extended functionality of rendering depth in colors. It works by splatting 3D Gaussians into the image plane via approximating the integral projection of the influence function $f(\cdot)$ along the depth dimension in pixel coordinates. The center of the Gaussian $\mu$, radius $r$, and depth $d$ (in camera coordinates) is splatted using the standard point rendering formula:
@@
where $K$ is the camera intrinsic matrix, $E_{t}$ is the extrinsic matrix capturing the rotation and translation of the camera at frame $t$, $l$ is the focal length. The influence of all Gaussians on this pixel can be combined by sorting the Gaussians in depth order and performing front-to-back volume rendering using the Max volume rendering formula [23]:
@@
where $d_{i}$ denotes the depth of each Gaussian. By setting $d_{i}=1$, we can calculate a silhouette, $Sil_{\rm pix}=D_{\rm pix}(d_{i}=1)$, which assists in determining whether a pixel is visible in the current view [16]. This aspect of visibility is essential for camera pose estimation, as it relies on the current reconstructed map. Additionally, it is also employed in map reconstruction, where new Gaussians are introduced in pixels lacking sufficient information.
@@ 3.2.3 Map Reconstruction
$\mathcal{L}_{\rm mapping}=\mathcal{U}_{t}\sum_{\rm pix}\lambda_{D}|D_{\rm pix}^{GT}-D_{\rm pix}|+\lambda_{C}\mathcal{L}_{C}+\lambda_{S}\mathcal{L}_{S}$ (8)
where $\mathcal{L}_{C}$ and $\mathcal{L}_{S}$ are weighted SSIM loss [17] with respect to appearance image and semantic image:
@@
$\lambda_{D}$, $\lambda_{C}$, $\lambda_{S}$, and $\alpha$ are predefined hyperparameters, and $\mathcal{U}_{t}$ is the uncertainty score defined in Sec. 3.2.2.
Compared to existing NeRF-based approaches [47, 15, 19, 46] that necessitate complex model architectures and feature fusion strategies, SGS-SLAM adopts explicit Gaussian representation for mapping. This approach allows for direct gradient flow to each parameter, resulting in high rendering speeds and optimal reconstruction quality. Compared to recent Gaussian-based methods [16, 40], SGS-SLAM incorporates geometric, appearance, and semantic features for multi-channel rendering. This enables the joint optimization of parameters across different channels, remarkably enhancing the efficiency and effectiveness of both mapping and segmentation processes.
@@ 4.1.1 Datasets
We evaluate our method on both synthetic and real-world datasets. To compare with other neural implicit SLAM methods, we evaluate synthetic scenes from Replica dataset [34] and real-world scenes from ScanNet [4] and ScanNet++ [43] datasets. The ground-truth camera pose and semantic map of Replica are offered from simulation, and the ground-truth camera pose of ScanNet is generated by BundleFusion [5]. The ground-truth 2D semantic label is provided by the dataset.
@@ 4.1.3 Baselines
We compare the tracking and mapping with state-of-the-art methods iMap [35], NICE-SLAM [47], Co-SLAM [37], ESLAM [15], and SplaTAM [16]. For semantic segmentation accuracy, we compare with NIDS-SLAM [11], DNS-SLAM [19], and SNI-SLAM [46].
@@ 4.2 Evaluation of Mapping and Localization
We show quantitative measures of reconstruction quality using the Replica dataset [34] in Tab. 1. Our method demonstrates state-of-the-art performance. Compared to other baseline methods, our approach attains notably superior outcomes, outperforming them by a margin of 10dB in PSNR.
Figure 2: Qualitative comparison of our method and the baselines for reconstruction across three scenes from the Replica Dataset [34], with key details accentuated using colorful boxes. The results demonstrate that our method delivers more high-fidelity and robust reconstructions, particularly by capturing more detailed features of the objects in the scene.
In Fig. 2, we present the reconstruction results of three chosen scenes, where regions of interest are accentuated with boxes in various colors. Our method exhibits high-fidelity reconstruction outcomes. Specifically, for small, intricately textured objects like a clock, socket, books on a tea table, and a lamp, our approach shows remarkable accuracy over NeRF-based methods. This is because Gaussians are capable of representing objects with complex textures and surfaces. Furthermore, NeRF-based methods often struggle with the over-smoothing issue, resulting in blurred edges on objects. In contrast, by utilizing an explicit Gaussian representation, SGS-SLAM precisely captures objects with clear edges, irrespective of their sizes. Compared with SplaTAM [16], which is also a Gaussian-based model, our approach utilizes semantic information for discerning object categories, recognizing visual appearance to determine texture, and applying geometric constraints to preserve accurate shapes. This combination enables our method to achieve thorough modeling of both objects and their surrounding environment. The combination of these constraints allows SGS-SLAM to capture fine-grained details of objects, offering high-fidelity and accurate reconstruction.
Table 1: Quantitative comparison of our method and the baselines in training view rendering on the Replica dataset [34]. Our method demonstrates SOTA performances in most cases among three metrics.
  Methods Metrics Avg.
Room0
Room1
Room2
Office0
Office1
Office2
Office3
Office4
[…]
24.42
22.12
22.47
24.52
29.07
30.34
19.66
22.23
24.94
NICE-SLAM
[…]
0.809
0.689
0.757
0.814
0.874
0.886
0.797
0.801
0.856
[…]
0.233
0.330
0.271
0.208
0.229
0.181
0.235
0.209
0.198
[…]
30.24
27.27
28.45
29.06
34.14
34.87
28.43
28.76
30.91
Co-SLAM
[…]
0.939
0.910
0.909
0.932
0.961
0.969
0.938
0.941
0.955
[…]
0.252
0.324
0.294
0.266
0.209
0.196
0.258
0.229
0.236
[…]
29.08
25.32
27.77
29.08
33.71
30.20
28.09
28.77
29.71
ESLAM
[…]
0.929
0.875
0.902
0.932
0.960
0.923
0.943
0.948
0.945
[…]
0.336
0.313
0.298
0.248
0.184
0.228
0.241
0.196
0.204
[…]
33.98
32.48
33.72
34.96
38.34
39.04
31.90
29.70
31.68
SplaTAM
[…]
0.969
0.975
0.970
0.982
0.982
0.982
0.965
0.950
0.946
[…]
0.099
0.072
0.096
0.074
0.083
0.093
0.100
0.118
0.155
[…]
34.66 32.50 34.25 35.10 38.54 39.20 32.90 32.05 32.75
Ours
[…]
0.973 0.976 0.978
0.981
0.984
0.980
0.967 0.966 0.949
[…]
0.096 0.070 0.094 0.070
0.086
0.087
0.101
0.115 0.148
Table 2: Quantitative comparison in terms of Depth L1, ATE, memory usage, and FPS between our method and the baselines on the Replica dataset [34]. The values represent the average outcomes across eight scenes. The results of baselines are retrieved from [46]. Our method remarkably outperforms the baselines at Depth and ATE evaluations, and performs fairly on FPS metrics. Note that for parameter size, we only count the parameters of the model. SOTA performances are highlighted.
  Methods
Depth L1 [cm]$\downarrow$
ATE Mean [cm]$\downarrow$
ATE RMSE [cm]$\downarrow$
@@ Methods
iMAP
4.645
3.118
4.153
9.92
2.23
1.82
NICE-SLAM
1.903
1.795
2.503
13.70
0.20
0.20
Co-SLAM
1.513
0.935
1.059
17.24
10.20 6.41
ESLAM
1.180
0.520
0.630
18.11
3.62
3.02
Ours 0.356 0.327 0.412
5.27
3.52
2.11
Tab. 2 displays the tracking evaluation results on the Replica dataset [34]. Our method excels in achieving the highest level of depth L1 loss (cm) and minimal ATE error, surpassing baseline methods by 70% in terms of depth loss and 34% in terms of ATE RMSE (cm). This exceptional performance can be attributed to our precise scene reconstruction, which provides finely-detailed rendering results. The high-quality rendering, in turn, contributes to accurate camera pose estimation based on the established map by preventing incorrect geometric reconstruction, which could otherwise result in inaccurate tracking outcomes. Additionally, utilizing features from different channels of Gaussians, such as geometry, appearance, and semantic information, provides multiple levels of supervision, resulting in a more robust and accurate tracking capability.